OpenStudy (anonymous):

evaluate g(n-5) if g(x)= x^2-6 / 2x

OpenStudy (anonymous):

Please help me!

hero (hero):

Replace x with (n-5) in g(x): \[g(x) = \frac{x^2 - 6}{2x}\]\[g(n - 5) = \frac{(n-5)^2 - 6}{2(n-5)}\] Now you have to simplify the expression for \(g(n-5)\)

hero (hero):

Do you plan to simplify it bro? Or are you just going to post answer choices?

OpenStudy (anonymous):

But how you simplify it?